OpenStudy (anonymous):

What is the first step in solving the equation 5x − 3 = 22? Add 3 to both sides of the equation. Divide both sides of the equation by −5. Divide both sides of the equation by −3. Subtract 3 from both sides of the equation.

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (anonymous):

since that would isolate 5x

OpenStudy (anonymous):

5x - 3 = 22 add 3 to both sides 5x = 25 divide by 5 5x/5 = 25/5 5/5 cancel out so x= 25/5